# J-VAR: Analysis of RR Lyrae light curves in seven optical bands

### v1

@@ 4.3 Residuals
Residuals were defined as the differences between data and fits, and the ratio of residuals to magnitude errors represents the units of error by which the fit differs from the data. This analysis is crucial for assessing the quality of the fit. Figure 6 shows the distribution of residuals normalized by magnitude errors across the seven optical bands in J-VAR DR1, with a Gaussian fit overlaid for comparison. The mean and standard deviation of the Gaussian were estimated separately for each filter and are presented in Table 4.
@@
The estimation of empirical light curves from J-VAR data or the use of different libraries (e.g. Baeza-Villagra et al. 2025) is needed to get additional clues about the origin of the enhanced dispersion, highlight important areas for improvement in template libraries and error modeling in variable star photometry. This is beyond the scope of the present paper and would be addressed in future work.
@@ 5.1 Comparison of periods
A threshold of $\Delta P<0.01$ days was adopted to evaluate the agreement between the two datasets (Fig. 7). Across all bands, period agreement rates peak near $15$ mag, with RRab stars achieving roughly 8080% consistency and RRc stars performing even better at around 9090%. The agreement rate drops for fainter stars, with RRab stars gradually decreasing, while RRc stars decline more quickly. For brighter stars, the agreement rate drops sharply. This is partly due to saturation effects in stars observed during exceptionally clear, photometric nights, where very high signal-to-noise ratios can cause flux overestimation or non-linear detector response. Zero-agreement cases may result from sparse data or limitations in the SDSS multiband template library, which may not fit well across all seven optical bands in J-VAR. For a threshold of $0.01$ days, RRab stars generally exhibit higher agreement rates than RRc stars, suggesting that fundamental mode pulsators were more reliably modeled by the template fitting method.
The SDSS multiband template library provides a more comprehensive understanding of RR Lyrae stars by integrating data from various bands. This allows for improved classification, more accurate period determination, and better insights into pulsation processes, e.g., mass loss, binary interactions (Bobrick et al., 2024). Moreover, a recent study (Ngeow & Cheng, 2025) includes a template library for $y$-band in the existing SDSS ugrizugriz bands. This inclusion provides an extended wavelength range, especially at longer wavelengths, improves the overall fit of the light curves, and facilitates a better classification of RR Lyrae subclasses. This is considered essential for understanding stellar evolution within our galaxy, particularly in populations such as the Galactic halo and globular clusters. By determining the subtype RRab, RRc, or RRd, more accurate estimations of the age and metallicity of these systems can be made. The creation of additional template libraries for the J-VAR medium and narrow bands is beyond the scope of the present work, but the residual analysis in Sect. 4.3 suggests that the current broadband templates capture most of the features present in the data.
@@ 5.2 Amplitude trends in the Gaia and ZTF surveys
The normalized amplitude in the Gaia $G$-band, and the ZTF $g$ and $r$ bands were computed to compare with the observed J-VAR trend. To minimize selection effects, only the RR Lyrae stars in the J-VAR DR1 catalog were used. The amplitudes for ZTF were obtained from its DR3 RR Lyrae catalog, J/ApJS/249/18222https://vizier.cds.unistra.fr/viz-bin/VizieR-3?-source=J/ApJS/272/31. A pivot wavelength of $622.88$ nm for the Gaia $G$-band was assumed (Weiler, 2018), while the ZTF $g$ and $r$ bands (Masci et al., 2019) share similar profiles to the J-VAR ones.
The measured normalized amplitudes, shown in Fig.  5, were $\overline{A}\,(G)=1.08\pm 0.21$ for Gaia, and $\overline{A}\,(g)=1.38\pm 0.22$ and $\overline{A}\,(r)=0.98\pm 0.15$ for ZTF. The normalized amplitude of the Gaia $G$-band follows the J-VAR exponential trend; it lies slightly above the exponential curve. This deviation stems from the broad nature of the Gaia $G$-band, which spans both blue and red wavelengths. Since most RR Lyrae stars are intrinsically blue, the $G$-band, despite its redder effective wavelength, still captures significant flux from the blue spectrum. Moreover, Gaia $G$-band magnitudes derive from fluxes convolved over a wide spectral range, introducing a non-linear response compared to narrower bands. This broad convolution may enhance measured variability relative to what a narrow, red band would detect, causing the slightly elevated median amplitude (Jordi et al., 2010).
@@ 6 Conclusions
The Bailey diagrams constructed using J-VAR DR1 periods and amplitudes successfully recover the expected sequences for RRab and RRc stars in all seven optical bands in J-VAR. The known trend with wavelength is recovered, with bluer passbands having larger amplitudes than redder passbands (e.g. Catelan, 2009). This trend was quantified with an exponential function after normalization to the $r$-band amplitudes, providing a benchmark for theoretical models. The normalized amplitudes provided by Gaia and ZTF for common RR Lyrae stars are compatible with the J-VAR results.
@@
This study also sets the stage for exploring links between RR Lyrae and Blue Horizontal Branch stars, shedding light on shared evolutionary paths and the Milky Way’s formation history (Carretta et al., 2004). The methods developed here may be extended to RR Lyrae populations in other Local Group galaxies, enhancing models of stellar evolution.
@@
The J-VAR DR1 data used are publicly available at CEFCA’s catalog portal [^3] https://archive.cefca.es/catalogues/jvar-dr1. The derived parameters for the $315$ RR Lyrae studied in this work are available at the table ?? of the J-VAR DR1 database and at CDS XXX.
Acknowledgements.
S. K. has received financial support from the Aragonese Government through the Research Groups E16_23R and the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grant PID2021-124918NB-C41. A. E. acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. Based on observations made with the JAST80 telescope and T80Cam camera for the J-PLUS project at the Observatorio Astrofísico de Javalambre (OAJ), in Teruel, owned, managed, and operated by the Centro de Estudios de Física del Cosmos de Aragón (CEFCA). We acknowledge the OAJ Data Processing and Archiving Department (UPAD; Cristóbal-Hornillos et al. 2012) for reducing and calibrating the OAJ data used in this work, as well as for the publication of the data products through a dedicated web portal. Funding for OAJ, UPAD, and CEFCA has been provided by the Governments of Spain and Aragón through the Fondo de Inversiones de Teruel and their general budgets; the Aragonese Government through the Research Groups E96, E103, E16_17R, E16_20R and E16_23R; the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grants PID2021-124918NB-C41, PID2021-124918NB-C42, PID2021-124918NA-C43, and PID2021-124918NB-C44; the Spanish Ministry of Science, Innovation and Universities (MCIU/AEI/FEDER, UE) with grant PGC2018-097585-B-C21; the Spanish Ministry of Economy and Competitiveness (MINECO) under AYA2015-66211-C2-1-P, AYA2015-66211-C2-2, AYA2012-30789, and ICTS-2009-14; and European FEDER funding (FCDD10-4E-867, FCDD13-4E-2685). Any other J-VAR related acknowledgments.
@@ Appendix A J-VAR DR1 RR Lyrae Catalog
This appendix describes the J-VAR DR1 + Gaia DR3 catalog of light curve analyses for $315$ RR Lyrae stars. The catalog includes both observational data and derived pulsation parameters. Table 5 summarizes the number of sources with available periods and amplitudes in each of the seven J-VAR optical bands. Table 6 provides a detailed overview of the catalog contents.
@@
Table 6: Description of the J-VAR DR1 + Gaia DR3 catalog of light curve analysis of $315$ RR Lyrae.

### v2

@@ 4.3 Residuals
Residuals were defined as the differences between data and fits, and the ratio of residuals to magnitude errors represents the units of error by which the fit differs from the data. This analysis is crucial for assessing the quality of the fit. Figure 6 shows the distribution of residuals normalized by magnitude errors across the seven optical bands in J-VAR DR1, with a Gaussian fit overlaid for comparison. The mean and standard deviation of the Gaussian were estimated separately for each filter and are presented in Tab. 4.
@@
The estimation of empirical light curves from J-VAR data or the use of different libraries (e.g. baeza25) is needed to get additional clues about the origin of the enhanced dispersion, highlighting important areas for improvement in template libraries and error modeling in variable star photometry. This is beyond the scope of the present paper and will be addressed in future work.
@@ 5.1 Comparison of periods
A threshold of $\Delta P<0.01$ days was adopted to evaluate the agreement between the two datasets (Fig. 7). The figure presents the fraction of J-VAR periods that agree with the Gaia values within our adopted threshold, shown as a function of the Gaia $G$ magnitude. Across all bands, period agreement rates peak near $15$ mag, with RRab stars achieving roughly 8080% consistency and RRc stars performing even better at around 9090%. The agreement rate drops for fainter stars, with RRab stars gradually decreasing, while RRc stars decline more quickly. For brighter stars, the agreement rate drops sharply. This is partly due to saturation effects in stars observed during exceptionally clear, photometric nights, where very high signal-to-noise ratios can cause non-linear detector response. Zero-agreement cases may result from sparse data or limitations in the SDSS multiband template library, which may not fit well across all seven optical bands in J-VAR. For a threshold of $0.01$ days, RRab stars exhibit higher agreement rates than RRc stars, suggesting that fundamental mode pulsators were more reliably modeled by the template fitting method.
The SDSS multiband template library provides a more comprehensive understanding of RR Lyrae stars by integrating data from various bands. This allows for improved classification, more accurate period determination, and better insights into pulsation processes, e.g., mass loss, binary interactions (Bobrick_2024). Moreover, a recent study (Ngeow_2025) includes a template library for $y$-band in the existing SDSS ugrizugriz bands. This inclusion provides an extended wavelength range, especially at longer wavelengths, improves the overall fit of the light curves, and facilitates a better classification of RR Lyrae subclasses. This is considered essential for understanding stellar evolution within our galaxy, particularly in populations such as the Galactic halo and globular clusters. By determining the subtype RRab, RRc, or RRd, more accurate estimations of the age and metallicity of these systems can be made. The creation of additional template libraries for the J-VAR medium and narrow bands is beyond the scope of the present work, but the residual analysis in Sect. 4.3 suggests that the current broadband templates capture most of the features present in the data.
@@ 5.2 Amplitude trends in the Gaia and ZTF surveys
The normalized amplitude in the Gaia $G$-band, and the ZTF $g$ and $r$ bands were computed to compare with the observed J-VAR trend. To minimize selection effects, only the RR Lyrae stars in the J-VAR DR1 catalog were used. A pivot wavelength of $622.88\penalty 10000\ \text{nm}$ for the Gaia $G$-band was assumed (Weller2018), while the ZTF $g$ and $r$ bands share profiles similar to those of J-VAR.
The measured normalized amplitudes, shown in Fig. 5, were $\overline{A}\,(G)=1.08\pm 0.21$ for Gaia, and $\overline{A}\,(g)=1.38\pm 0.22$ and $\overline{A}\,(r)=0.98\pm 0.15$ for ZTF. The normalized amplitude of the Gaia $G$-band follows the J-VAR exponential trend; it lies slightly above the exponential curve. This deviation stems from the broad nature of the Gaia $G$-band, which spans both blue and red wavelengths. Since most RR Lyrae stars are intrinsically blue, the $G$-band, despite its redder effective wavelength, still captures significant flux from the blue spectrum. Moreover, Gaia $G$-band magnitudes derive from fluxes convolved over a wide spectral range, introducing a non-linear response compared to narrower bands. This broad convolution may enhance measured variability relative to what a narrow, red band would detect, causing the slightly elevated median amplitude (Jordi2010).
@@ 6 Conclusions
The Bailey diagrams constructed using J-VAR DR1 periods and amplitudes successfully recover the expected sequences for RRab and RRc stars in all seven optical bands in J-VAR. The known trend with wavelength is recovered, with bluer passbands having larger amplitudes than redder passbands (e.g. Catelan2009). This trend was quantified with an exponential function after normalization to the $r$-band amplitudes, providing a benchmark for theoretical models. The normalized amplitudes provided by Gaia and ZTF for common RR Lyrae stars are compatible with the J-VAR results.
@@
This study also sets the stage for exploring links between RR Lyrae and Blue Horizontal Branch stars, shedding light on shared evolutionary paths and the Milky Way’s formation history (Carretta2004). The methods developed here may be extended to RR Lyrae populations in other Local Group galaxies, enhancing models of stellar evolution.
Acknowledgements
S. K. has received financial support from the Aragonese Government through the Research Groups E16_23R and the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grant PID2021-124918NB-C41.
A. E. acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA.
F. J-E. acknowledges that this research has been partially funded by MICIU/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.
Based on observations made with the JAST80 telescope and T80Cam camera for the J-PLUS project at the Observatorio Astrofísico de Javalambre (OAJ), in Teruel, owned, managed, and operated by the Centro de Estudios de Física del Cosmos de Aragón (CEFCA). We acknowledge the OAJ Data Processing and Archiving Department (UPAD; upad) for reducing and calibrating the OAJ data used in this work, as well as for the publication of the data products through a dedicated web portal.
Funding for OAJ, UPAD, and CEFCA has been provided by the Governments of Spain and Aragón through the Fondo de Inversiones de Teruel and their general budgets; the Aragonese Government through the Research Groups E96, E103, E16_17R, E16_20R and E16_23R; the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grants PID2021-124918NB-C41, PID2021-124918NB-C42, PID2021-124918NA-C43, and PID2021-124918NB-C44; the Spanish Ministry of Science, Innovation and Universities (MCIU/AEI/FEDER, UE) with grant PGC2018-097585-B-C21; the Spanish Ministry of Economy and Competitiveness (MINECO) under AYA2015-66211-C2-1-P, AYA2015-66211-C2-2, AYA2012-30789, and ICTS-2009-14; and European FEDER funding (FCDD10-4E-867, FCDD13-4E-2685).
We thank the J-VAR Collaboration, as well as Álvaro Álvarez-Candal and Vinicius Placco of the J-PLUS Collaboration, for their valuable comments.
@@ Acknowledgements
The J-VAR DR1 data used are publicly available at CEFCA’s catalog portal [^4] https://archive.cefca.es/catalogues/jvar-dr1. The derived parameters for the $315$ RR Lyrae studied in this work are available at Zenodo [^5] https://zenodo.org/records/17601696.
@@ Appendix A J-VAR DR1 RR Lyrae Catalog
This appendix describes the J-VAR DR1 + Gaia DR3 catalog of light curve analyses for $315$ RR Lyrae stars. The catalog includes both observational data and derived pulsation parameters. Tab. 5 summarizes the number of sources with available periods and amplitudes in each of the seven J-VAR optical bands. Tab. 6 provides a detailed overview of the catalog contents.
@@
Table 6: Description of the J-VAR DR1 ++ Gaia DR3 catalog of light curve analysis of $315$ RR Lyrae.